# End-to-End and Highly-Efficient Differentiable Simulation for Robotics

### v1

@@ Abstract
Over the past few years, robotics simulators have largely improved in efficiency and scalability, enabling them to generate years of simulated data in a few hours. Yet, efficiently and accurately computing the simulation derivatives remains an open challenge, with potentially high gains on the convergence speed of reinforcement learning and trajectory optimization algorithms, especially for problems involving physical contact interactions. This paper contributes to this objective by introducing a unified and efficient algorithmic solution for computing the analytical derivatives of robotic simulators. The approach considers both the collision and frictional stages, accounting for their intrinsic nonsmoothness and also exploiting the sparsity induced by the underlying multibody systems. These derivatives have been implemented in C++, and the code will be open-sourced in the Simple simulator. They depict state-of-the-art timings ranging from 5 us for a 7-dof manipulator up to 95 us for 36-dof humanoid, outperforming alternative solutions by a factor of at least 100. For videos additional details, please see our project website.
**footnotetext: Equal contribution.
Keywords: Differentiable physics, Differentiable optimization, Simulation, Nonsmooth dynamics, Model-based learning
1 Introduction
@@
Contributions. In this paper, we present a comprehensive framework for differentiable simulation that combines differentiable rigid-body dynamics, differentiable collision detection, and differentiable contact resolution. We notably introduce an implicit differentiation scheme to compute the gradients of the NCP associated with the frictional contact problem without any relaxation and chain it with rigid body dynamics algorithms to finely exploit the kinematic sparsity of the problem. Our approach achieves substantial computational speedups, with gradient computations up to 100 times faster than current state-of-the-art methods in robotics, while avoiding any physical relaxation or geometrical approximation on meshes. We validate the effectiveness of our method by applying it to complex inverse problems, including the estimation of initial conditions and inverse dynamics through contact. To support reproducibility and further research, we will make our code publicly available after the review process.
Paper organization. In Sec. 2, we provide a background on collision detection, frictional dynamics, rigid multi-body dynamics, and implicit differentiation techniques. Sec. 3 corresponds to the core contribution of this paper. We specify the computational graph of a physics engine and explain how the combination of gradients of rigid-body dynamics, collision detection, and contact forces make simulation end-to-end differentiable. Additionally, we show how implicit differentiation and rigid body algorithms can be leveraged to compute the derivatives of multibody frictional contact problems efficiently, including collision geometry contributions. In Sec. 4, the efficiency of our approach is benchmarked on several advanced robotics systems, and we leverage our differentiable physics engine to tackle various estimation and control problems. Sec. 5 discusses this work’s limitations and how it could set the stage for future developments in model-based approaches for robotics.
Table 1: Differentiable physics engines for robotics.
Physics engine Contact Model $\partial$(Contacts) $\partial$(Collisions) MuJoCo MJX [2] CCP Auto-diff meshes + primitives Nimble [10], [8] LCP Implicit meshes + primitives Dojo[13] NCP Implicit (not considered) Ours NCP Implicit meshes + primitives
@@
2 Background
This section reviews the two fundamental stages of modern robotic simulators: collision detection, contact modelling and multibody dynamics. We also review the notion of implicit differentiation, which is at the core of our approach.
2.1 Collision detection
@@
2.2 Frictional contact dynamics
@@
where $G$ is the so-called Delassus matrix [23] that gives the system inverse inertia projected on the contacts. It is a linear operator mapping contact forces to contact velocities. $\bm{g}$ is the free velocity of the contact. $\mathcal{K}_{\mu}$ is a second-order cone with aperture angle $\operatorname{atan}(\mu)$, $\mathcal{K}_{\mu}^{*}=\mathcal{K}_{1/\mu}$ its dual cone and $\Gamma_{\mu}(\bm{\sigma})=[0,0,\mu\|\bm{\sigma}_{T}\|]$ is the so-called De Saxcé correction [24, 25] enforcing the Signorini condition [25, 16]. (1) can be solved by interior point methods [13], projected Gauss-Seidel [26] or ADMM-based approaches [25, 27, 28].
2.3 Multibody frictional contact dynamics
@@
where $\bm{\dot{v}}_{f}=M^{-1}\left(\bm{\tau}-b\right)$ is the free acceleration term and $\Delta t$ is the time step. The acceleration term $\bm{\dot{v}}_{f}+M^{-1}J_{c}^{\top}\bm{\lambda}$ of (3) can be efficiently computed with the Articulated Body Algorithm (ABA) [30]. Next, we use the shorthand $\text{ABA}(\bm{q},\bm{v},\bm{\tau},\bm{\lambda})=\bm{\dot{v}}_{f}+M^{-1}J_{c}^{\top}\bm{\lambda}$.
Multiplying (3) by $J_{c}$, we recover from the definition of the Jacobian $J_{c}\bm{v}^{+}=\bm{\sigma}=G\bm{\lambda}+g$ associated to the NCP (1) in the case of multibody dynamics. This yields the expression of the Delassus matrix $G=J_{c}M^{-1}J_{c}^{\top}$ and the free contact velocity vector $g=J_{c}(\bm{v}+\Delta t\bm{\dot{v}}_{f})$. For poly-articulated rigid-body systems, $G$ depends on $\bm{q}$, and $g$ depends on $\bm{q},\bm{v},\bm{\tau}$. In this respect, the associated NCP is conditioned by $\bm{q},\bm{v},\bm{\tau}$. Note also that the contact Jacobian $J_{c}$ depends on $\bm{q}$, first through the kinematic structure of the system and second through the contact frame $c(\bm{q})$.
2.4 Implicit differentiation
@@
This approach has been successfully applied to differentiate Quadratic Programming (QP) problems in [31] and generalized to convex cone programs [32] and LCPs [8]. More generally, it allows incorporating optimization layers in the differentiable programming paradigm. In Sec. 3.2, we extend this approach to the NCP case and propose a method to compute the gradients of the contact forces efficiently.
3 Efficient differentiable simulation
This section details the core contribution of this paper, namely a comprehensive framework for differentiable simulation that combines differentiable rigid-body dynamics, differentiable collision detection, and differentiable contact resolution We show the link between the derivatives of the multibody dynamics, the frictional contact dynamics, and the derivatives of collision detection. We introduce an efficient algorithmic solution to compute the derivatives associated with the contact NCP by solving a reduced system of equations of minimal dimension resulting from its implicit differentiation.
3.1 Chaining rigid-body dynamics derivatives and NCP derivatives
From Sec. 2, the simulation equations (3) can be restated using (ABA) and the solution to the (NCP) problem (1) as
@@ 3.1 Chaining rigid-body dynamics derivatives and NCP derivatives
$\bm{v}^{+}$ $=\bm{v}+\Delta t\,\text{ABA}(\bm{q},\bm{v},\bm{\tau},\bm{\lambda}^{*}).$ (5)
Next, we consider forward-mode differentiation setup [33], i.e., we aim to compute the Jacobian $\frac{\text{d}v^{+}}{\text{d}\theta}$ where $\theta\in\mathbb{R}^{n_{p}}$ represents any subset of the inputs $\{\bm{q},\bm{v},\bm{\tau}\}$ or physical parameters. Our approach can be efficiently adapted to the reverse mode by applying the computational trick introduced in [31].
Differentiating (5) leads to
$\frac{\text{d}\bm{v}^{+}}{\text{d}\theta}$ $=\underbrace{\frac{\text{d}\bm{v}}{\text{d}\theta}+\Delta t\left(\frac{\partial\text{ABA}}{\partial\bm{q}}\frac{\text{d}\bm{q}}{\text{d}\theta}+\frac{\partial\text{ABA}}{\partial\bm{v}}\frac{\text{d}\bm{v}}{\text{d}\theta}+\frac{\partial\text{ABA}}{\partial\bm{\tau}}\frac{\text{d}\bm{\tau}}{\text{d}\theta}\right)}_{\left.\frac{\text{d}\bm{v}^{+}}{\text{d}\theta}\right|_{\lambda=\lambda^{*}}}+\Delta t\frac{\partial\text{ABA}}{\partial\bm{\lambda}}\frac{\text{d}\bm{\lambda}^{*}}{\text{d}\theta},$ (6)
where we identify the term $\left.\frac{\text{d}\bm{v}^{+}}{\text{d}\theta}\right|_{\lambda=\lambda^{*}}$ of derivatives, considering that $\bm{\lambda}^{*}$ does not vary. The ABA derivatives $\frac{\partial\text{ABA}}{\partial\bm{q},\bm{v},\bm{\tau}}$ and $\frac{\partial\text{ABA}}{\partial\bm{\lambda}}=M^{-1}(\bm{q})J_{c}^{\top}(\bm{q})$ can be efficiently computed via rigid-body algorithms [34] and are, for instance, available in Pinocchio [35]. At this stage, it is worth noting that $\frac{\partial\text{ABA}}{\partial\bm{q}}$ also depends on the geometry of the contact through the contact Jacobian $J_{c}(\bm{q},c(\bm{q}))$. The classical ABA rigid body algorithm gives the derivatives related to the first variable, and we need to add a term related to the variations of $J_{c}$ induced by the variation of the contact point $c(\bm{q})$ (see Section 3.4).
@@
3.2 Implicit differentiation of the NCP
The dynamics induced by the NCP (1) is inherently nonsmooth as it can switch on three modes. These modes correspond to the active set of (1) and result in different gradients for the contact dynamics. Our approach considers scenarios with multiple contact points and requires to identify the mode for each contact. For clarity purposes, we present the equations for a single contact point in the case of each of the three modes.
@@ 3.2 Implicit differentiation of the NCP
Mode 3 - Sliding contact (sld). When evolving in this regime, the contact point moves on the contact surface, which implies a null normal velocity ($\bm{\sigma}_{N}=0$) and a non-null tangential velocity ($\|\bm{\sigma}_{T}\|>0$). Moreover, from the MDP, tangential contact forces should lie on the boundary of the cone and in the opposite direction of the tangential velocity ($\bm{\lambda}^{*}_{T}=-\mu\bm{\lambda}^{*}_{N}\frac{\bm{\sigma}_{T}}{\|\bm{\sigma}_{T}\|}$). This additionally implies that $\frac{\mathrm{d}\bm{\lambda}^{*}}{\text{d}\theta}$ should be in the plane tangent to the friction cone as illustrated in Fig. 1 and allows reducing the search space to a 2D plane via a simple change of variable $\frac{\mathrm{d}\bm{\lambda}^{*}}{\text{d}\theta}=R\frac{\mathrm{d}\bm{\tilde{\lambda}}}{\text{d}\theta}$ where $R=\begin{pmatrix}\frac{\bm{\lambda}}{\|\bm{\lambda}\|}&\bm{e}_{z}\times\frac{\bm{\sigma}_{T}}{\|\bm{\sigma}_{T}\|}\end{pmatrix}\in\mathbb{R}^{3\times 2}$. Therefore, differentiating these equations yields the following conditions on the gradients
@@
where $P=\begin{pmatrix}H(\bm{\sigma}_{T})&0_{2\times 1}\\ 0_{1\times 2}&1\end{pmatrix}\in\mathbb{R}^{3\times 3}$, $H(x)=\frac{1}{\alpha}\left(\text{Id}-\frac{x}{\|x\|}\frac{x}{\|x\|}^{\top}\right)\in\mathbb{R}^{2\times 2}$, $\tilde{G}={R}^{\top}PG{R}+Q$ with $Q=\begin{pmatrix}0&0\\ 0&1\end{pmatrix}\in\mathbb{R}^{2\times 2}$ and $\alpha=\frac{\|\bm{\sigma}_{T}\|}{\mu\bm{\lambda}_{N}}$. We refer to the appendix for the detailed derivation.
3.3 Efficient computation: exploiting kinematic-induced sparsity
One should first identify the active contact modes to obtain the equations of all contact points. We denote $\mathcal{A}_{\text{brk}}$, $\mathcal{A}_{\text{stk}}$, and $\mathcal{A}_{\text{sld}}$ as the sets of contact indices corresponding to the breaking, sticking, and sliding contacts respectively. The dynamics of the different contacts are coupled through the Delassus matrix $G$. Thus, we construct a matrix $A\in\mathbb{R}^{(3n_{stk}+2n_{sld})\times(3n_{stk}+2n_{sld})}$ where from $G$ we remove the blocks related to $\mathcal{A}_{\text{brk}}$ and modify the lines and columns of $G$ related to $\mathcal{A}_{\text{sld}}$ by following the pattern of $\tilde{G}$ presented in (20). Once this is done, the reduced linear system on $X$, the stacking of $\text{d}\bm{\lambda}^{*}$ and $\text{d}\tilde{\bm{\lambda}}$, is obtained by concatenating the corresponding right-hand side of (8) and (20). We obtain the linear system corresponding to implicit differentiation

### v2

# Highly-Efficient Differentiable Simulation for Robotics

@@ Abstract
Over the past few years, robotics simulators have significantly improved in computational speed and scalability, enabling them to generate years of simulated data for complex robotic systems in a matter of minutes or hours. However, despite these advancements, the efficient and accurate computation of simulation derivatives remains an open challenge. Addressing this challenge could substantially accelerate the convergence of reinforcement learning and trajectory optimization algorithms, particularly for contact-rich problems. This paper addresses this challenge by introducing a unifying framework for robotic simulation that comprehensively considers all the factors in simulation, including dynamics, collisions, and friction. It results in an efficient algorithm that leverages implicit differentiation to compute the analytical derivatives of the simulation. It explicitly accounts for the intrinsic non-smoothness of collision and frictional simulation stages while exploiting the sparsity in dynamics induced by the multi-body system structure. These derivatives have been implemented in C++, and the code will be open-sourced after the review process to facilitate broader applications in robotics, such as simulation-driven learning or real-time control. Benchmark results demonstrate state-of-the-art performance, with timings ranging from 5 us for a 7-dof manipulator to 95 us for a 36-dof humanoid—an improvement of at least two orders of magnitude over alternative methods, such as automatic differentiation.
11footnotetext: Equal contribution.
I Introduction
I-A Context
@@
I-B Contributions
In this paper, we present a comprehensive framework for differentiable simulation that combines differentiable rigid-body dynamics, differentiable collision detection, and differentiable contact resolution. We notably introduce an implicit differentiation scheme to compute the gradients of the NCP associated with the frictional contact problem without any relaxation and chain it with rigid body dynamics algorithms to finely exploit the kinematic sparsity of the problem. Our approach achieves substantial computational speedups, with gradient computations up to 100 times faster than current state-of-the-art methods in robotics, while avoiding any physical relaxation or geometrical approximation on meshes. We validate the effectiveness of our method by applying it to complex inverse problems, including the estimation of initial conditions and inverse dynamics through contact. We also show that our gradients can be used in a policy learning context to improve the sample efficiency during training. To support reproducibility and further research, we will make our code publicly available after the review process.
I-C Paper organization
In Sec. II, we provide a background on collision detection, frictional dynamics, rigid multi-body dynamics, and implicit differentiation techniques. Sec. III corresponds to the core contribution of this paper. We specify the computational graph of a physics engine and explain how the combination of gradients of rigid-body dynamics, collision detection, and contact forces make simulation end-to-end differentiable. Additionally, we show how implicit differentiation and rigid body algorithms can be leveraged to compute the derivatives of multibody frictional contact problems efficiently, including collision geometry contributions. In Sec. IV, the efficiency of our approach is benchmarked on several advanced robotics systems. We also leverage our differentiable physics engine to tackle various estimation and control problems and to efficiently learn policies. Sec. V discusses this work’s limitations and how it could set the stage for future developments in model-based approaches for robotics.
TABLE I: Differentiable physics engines for robotics.
@@ I-C Paper organization
[…]
II Background
This section reviews the three fundamental stages of modern robotic simulators: collision detection, contact modeling and multibody dynamics. We also review the notion of implicit differentiation, which is at the core of our approach.
II-A Collision detection
@@
II-B Frictional contact dynamics
@@
where $G$ is the so-called Delassus matrix [12] that gives the system inverse inertia projected on the contacts. It is a linear operator mapping contact forces to contact velocities. $\bm{g}$ is the free velocity of the contact. $\mathcal{K}_{\mu}$ is a second-order cone with aperture angle $\operatorname{atan}(\mu)$, $\mathcal{K}_{\mu}^{*}=\mathcal{K}_{1/\mu}$ its dual cone and $\Gamma_{\mu}(\bm{\sigma})=[0,0,\mu\|\bm{\sigma}_{T}\|]$ is the so-called De Saxcé correction [1, 11] enforcing the Signorini condition [1, 30]. Problem (1) can be solved by interior point methods [24], projected Gauss-Seidel [27] or ADMM-based approaches [1, 8, 41].
II-C Multibody frictional contact dynamics
@@
where $\bm{\dot{v}}_{f}=M^{-1}\left(\bm{\tau}-b\right)$ is the free acceleration term and $\Delta t$ is the time step. The acceleration term $\bm{\dot{v}}_{f}+M^{-1}J_{c}^{\top}\bm{\lambda}$ of (3) correspond to the unconstrained forward dynamic (UFD) with exterior forces $\bm{\lambda}$ which can be efficiently computed with the Articulated Body Algorithm (ABA) [16]. In the remaining of the paper, we use the shorthand $\text{UFD}(\bm{q},\bm{v},\bm{\tau},\bm{\lambda})=\bm{\dot{v}}_{f}+M^{-1}J_{c}^{\top}\bm{\lambda}$. Multiplying Eq. (3) by $J_{c}$, we recover the contact velocity associated to the contact NCP (1) in the case of multibody dynamics:
$J_{c}\bm{v}^{+}=\bm{\sigma}=G\bm{\lambda}+g.$ (4)
This yields the expression of the Delassus matrix $G=J_{c}M^{-1}J_{c}^{\top}$ and the free contact velocity vector $g=J_{c}(\bm{v}+\Delta t\bm{\dot{v}}_{f})$. For poly-articulated rigid-body systems, $G$ depends on $\bm{q}$, and $g$ depends on $\bm{q},\bm{v},\bm{\tau}$. In this respect, the associated NCP is conditioned by $\bm{q},\bm{v},\bm{\tau}$. Note also that the contact Jacobian $J_{c}$ depends on $\bm{q}$, first through the kinematic structure of the system and second through the contact frame $c(\bm{q})$.
II-D Implicit differentiation
@@
This approach has been successfully applied to differentiate Quadratic Programming (QP) problems in [3] and generalized to convex cone programs [2] and LCPs [10]. More generally, it allows incorporating optimization layers in the differentiable programming paradigm. In Sec. III-B, we extend this approach to the NCP case and propose a method to compute the gradients of the contact forces efficiently.
III Efficient differentiable simulation
This section details the core contribution of this paper, namely a comprehensive framework for differentiable simulation that combines differentiable rigid-body dynamics, differentiable collision detection, and differentiable contact resolution. We show the link between the derivatives of multibody dynamics, frictional contact dynamics, and collision detection. We introduce an efficient algorithmic solution to compute the derivatives associated with the contact NCP by solving a reduced system of equations of minimal dimension resulting from its implicit differentiation.
III-A Chaining rigid-body dynamics derivatives and NCP derivatives
From Sec. II, the simulation equations (3) can be restated using unconstrained forward dynamic (UFD) and the solution to the (NCP) problem (1) as
@@ III-A Chaining rigid-body dynamics derivatives and NCP derivatives
$\bm{v}^{+}$ $=\bm{v}+\text{UFD}(\bm{q},\bm{v},\bm{\tau},\bm{\lambda}^{*})\Delta t.$ (6)
Next, we consider forward-mode differentiation setting [21], i.e., we aim to compute the Jacobian $\frac{\text{d}v^{+}}{\text{d}\theta}$ where $\theta\in\mathbb{R}^{n_{p}}$ represents any subset of the inputs $\{\bm{q},\bm{v},\bm{\tau}\}$ or physical parameters. Our approach can be efficiently adapted to the reverse mode by applying the computational trick introduced in [3]. Differentiating (6) leads to
$\frac{\text{d}\bm{v}^{+}}{\text{d}\theta}=$ $\frac{\partial\text{UFD}}{\partial\bm{\lambda}}\frac{\text{d}\bm{\lambda}^{*}}{\text{d}\theta}\Delta t$ (7)
$+\underbrace{\frac{\text{d}\bm{v}}{\text{d}\theta}+\left(\frac{\partial\text{UFD}}{\partial\bm{q}}\frac{\text{d}\bm{q}}{\text{d}\theta}+\frac{\partial\text{UFD}}{\partial\bm{v}}\frac{\text{d}\bm{v}}{\text{d}\theta}+\frac{\partial\text{UFD}}{\partial\bm{\tau}}\frac{\text{d}\bm{\tau}}{\text{d}\theta}\right)\Delta t}_{\left.\frac{\text{d}\bm{v}^{+}}{\text{d}\theta}\right|_{\lambda=\lambda^{*}}},$
where we identify the term $\left.\frac{\text{d}\bm{v}^{+}}{\text{d}\theta}\right|_{\lambda=\lambda^{*}}$ of derivatives, considering that $\bm{\lambda}^{*}$ does not vary. The derivatives $\frac{\partial\text{UFD}}{\partial\bm{q},\bm{v},\bm{\tau}}$ and $\frac{\partial\text{UFD}}{\partial\bm{\lambda}}=M^{-1}(\bm{q})J_{c}^{\top}(\bm{q})$ can be efficiently computed via rigid-body algorithms [6] and are, for instance, available in Pinocchio [7]. At this stage, it is worth noting that $\frac{\partial\text{UFD}}{\partial\bm{q}}$ also depends on the geometry of the contact through the contact Jacobian $J_{c}(\bm{q},c(\bm{q}))$. The classical ABA rigid body algorithm gives the derivatives related to the first variable, and we need to add a term related to the variations of $J_{c}$ induced by the variation of the contact point $c(\bm{q})$ (see Section III-D and Appendix C).
@@
III-B Implicit differentiation of the NCP
The dynamics induced by the NCP (1) is inherently nonsmooth as it can switch on three modes. These modes correspond to the active set of (1) and result in different gradients for the contact dynamics. Our approach considers scenarios with multiple contact points and requires the identification of the mode for each contact. For clarity purposes, we present the equations for a single contact point in the case of each of the three modes.
@@ III-B Implicit differentiation of the NCP
Mode 3 - Sliding contact (sld). In this regime, the contact point moves on the contact surface, which implies a null normal velocity ($\bm{\sigma}_{N}=0$) and a non-null tangential velocity ($\|\bm{\sigma}_{T}\|>0$). Moreover, from the MDP, tangential contact forces should lie on the boundary of the cone and in the opposite direction of the tangential velocity ($\bm{\lambda}^{*}_{T}=-\mu\bm{\lambda}^{*}_{N}\frac{\bm{\sigma}_{T}}{\|\bm{\sigma}_{T}\|}$). This additionally implies that $\frac{\mathrm{d}\bm{\lambda}^{*}}{\text{d}\theta}$ should be in the plane tangent to the friction cone as illustrated in Fig. 1 and allows reducing the search space to a 2D plane via a simple change of variable $\frac{\mathrm{d}\bm{\lambda}^{*}}{\text{d}\theta}=R\frac{\mathrm{d}\bm{\tilde{\lambda}}}{\text{d}\theta}$ where $R=\begin{pmatrix}\frac{\bm{\lambda}}{\|\bm{\lambda}\|}&\bm{e}_{z}\times\frac{\bm{\sigma}_{T}}{\|\bm{\sigma}_{T}\|}\end{pmatrix}\in\mathbb{R}^{3\times 2}$. Therefore, differentiating these equations yields the following conditions on the gradients
@@
where: $P=\begin{pmatrix}H(\bm{\sigma}_{T})&0_{2\times 1}\\ 0_{1\times 2}&1\end{pmatrix}\in\mathbb{R}^{3\times 3}$, $H(x)=\frac{1}{\alpha}\left(\text{Id}-\frac{x}{\|x\|}\frac{x}{\|x\|}^{\top}\right)\in\mathbb{R}^{2\times 2}$, $\tilde{G}={R}^{\top}PG{R}+Q$ with $Q=\begin{pmatrix}0&0\\ 0&1\end{pmatrix}\in\mathbb{R}^{2\times 2}$ and $\alpha=\frac{\|\bm{\sigma}_{T}\|}{\mu\bm{\lambda}_{N}}$. We refer to A for the detailed derivation.
III-C Efficient computation: exploiting kinematic-induced sparsity
First, one should identify the active contact modes to obtain the equations for all contact points. We denote $\mathcal{A}_{\text{brk}}$, $\mathcal{A}_{\text{stk}}$, and $\mathcal{A}_{\text{sld}}$ as the sets of contact indices corresponding to the breaking, sticking, and sliding contacts respectively. The dynamics of the different contacts are coupled through the Delassus matrix $G$. Thus, we construct a matrix $A\in\mathbb{R}^{(3n_{stk}+2n_{sld})\times(3n_{stk}+2n_{sld})}$ where from $G$ we remove the blocks related to $\mathcal{A}_{\text{brk}}$ and modify the lines and columns of $G$ related to $\mathcal{A}_{\text{sld}}$ by following the pattern of $\tilde{G}$ presented in (10). Once this is done, the reduced linear system on $X$, the stacking of $\text{d}\bm{\lambda}^{*}$ and $\text{d}\tilde{\bm{\lambda}}$, is obtained by concatenating the corresponding right-hand side of (9) and (10). We obtain the linear system corresponding to implicit differentiation